# Preconditioning without a preconditioner using randomized block Krylov subspace methods

### v2

@@ 4 Probabilistic bounds
In this section we prove probabilistic bounds for augmented block-CG. In light of Theorem 3.4, our strategy for deriving bounds for block-CG is simply to derive bounds for Nyström PCG. To do this, we follow the approach developed in [23] and use a deterministic bound for the condition number of the Nyström preconditioned system. The following is a minor generalization of Proposition 5.3 in [23] to allow arbitrary $\theta$. The proof is contained in Section C.1.
@@
Let $\mathbf{P}_{\mu}$ be the Nyström preconditioner Eq. 9 corresponding to the Nyström approximation $\mathbf{A}\langle\mathbf{K}\rangle$ for any $\mathbf{K}$ and shift parameter $\theta\geq 0$. Then
@@
In particular, Proposition 4.1 implies that if $\|\mathbf{A}-\mathbf{A}\langle\mathbf{K}\rangle\|\approx\lambda_{r+1}$ and $\theta\in[\lambda_{d},\lambda_{r+1}]$, then $\kappa(\mathbf{P}_{\mu}^{-1/2}\mathbf{A}_{\mu}\mathbf{P}_{\mu}^{-1/2})\approx\kappa_{r+1}(\mu)$.
@@
Our first result, which we prove in Section 4.1, is a “numerical analysis” style bound based on the error gurantees for Nyström low-rank approximation in [61].
@@
$s\geq\min\left\{2+\frac{\log(d)}{2},{\frac{3}{2}}+\frac{1}{4}\log\left(4+\frac{4r}{p-1}\sum_{i>r}\frac{\lambda_{i}^{2}}{\lambda_{r+1}^{2}}\right)\right\}.$
@@
Note that $\mathbf{A}\langle\mathbf{K}_{s}\rangle$ as defined in Eq. 12 can have rank as large as $\ell s$, so we might hope that roughly $s=r/\ell$ iterations suffice. Recent work [44, 8] implies this is more-or-less the case. In Section 4.2, we use this result to prove our second result, a “theoretical computer science” type bound.
@@
In contrast to Theorem 4.2, which requires $\ell\geq r+2$, Theorem 4.3 allows any choice of $\ell$. The bound reveals that we get exponential convergence at a rate roughly $1/\sqrt{\kappa_{r+1}(\mu)}$ after roughly $s\approx r/\ell$ matrix-loads (and hence $r\approx\ell s$ matrix-vector products). Note that the logarithmic dependence on eigenvalue gaps $\Delta$ is generally considered to be mild due to “smoothed-analysis” type effects of finite-precision arithmetic [44, 8].
@@
The fact that Theorems 4.2 and 4.3 guarantee an accurate result for all $\mu\geq 0$ with high probability (as opposed to a single value of $\mu$) will be important in our applications. In particular, it allows guarantees for computing the whole ridge-regression regularization path and will be necessary in the analysis of our algorithm for sampling Gaussian vectors.
@@
Bounds based on condition numbers (such as Theorem 4.2) are often pessimistic in practice. To determine how many iterations to run, it is more common to use some sort of a posteriori error estimate. For example, observing the residual $\|\mathbf{b}-\mathbf{A}_{\mu}\mathsf{bcg}_{t}^{(1)}(\mu)\|$ gives some indication of the quality of the solution. More advanced techniques can also be used [43].
@@ 4.1 Explicit bound
In this section, we prove Theorem 4.2. We do not attempt to optimize constants, opting instead for simple arguments and clean theorem statements.
@@
Suppose $\mathbf{\Omega}\in\mathbb{R}^{d\times(r+p)}$ is a random Gaussian matrix and define $\mathbf{K}_{s}:=[\mathbf{\Omega}~\mathbf{A}\mathbf{\Omega}~\cdots~\mathbf{A}^{s-1}\mathbf{\Omega}]$. Let $\mathbf{P}_{\mu}$ be the Nyström preconditioner Eq. 9 corresponding to the Nyström approximation $\mathbf{A}\langle\mathbf{K}_{s}\rangle$ for any shift parameter $\theta\in[\lambda_{d},\lambda_{r+1}]$. Suppose $p\geq 2$ and
$s\geq\min\left\{2+\frac{\log(d)}{2},\frac{3}{2}+\frac{1}{4}\log\left(4+\frac{4r}{p-1}\sum_{i>r}\frac{\lambda_{i}^{2}}{\lambda_{r+1}^{2}}\right)\right\}.$
@@
The proof is a simple consequence of Theorems 4.6 and 4.1. We begin by bounding $\|\mathbf{A}-\mathbf{A}\langle\mathbf{K}\rangle\|$. We first note that the minimum is always attained by the second term. Indeed, since $r\leq d$ and $\lambda_{i}/\lambda_{r+1}\leq 1$ for $i>r$, by setting $p=2$, we see $4r/(p-1)\sum_{i>r}\lambda_{i}^{2}/\lambda_{r+1}^{2}\leq 4d^{2}$. Now, using properties of the logarithm, $\log(4+4d^{2})=\log(1+d^{2})+\log(4)\leq\log(d^{2})+\log(2)+\log(4)=\log(8)+2\log(d)$. The claim then follows since $3/2+\log(8)/4\leq 2$.
With the condition on $s$, Theorem 4.6 guarantees that
(25) $\operatorname{\mathbb{E}}\Bigl{[}\|\mathbf{A}-\mathbf{A}\langle\mathbf{K}\rangle\|^{2}\Bigr{]}\leq\mathrm{e}^{1/2}\cdot\lambda_{r+1}^{2}.$
@@
(26) $\operatorname{\mathbb{P}}\Bigl{[}\|\mathbf{A}-\mathbf{A}\langle\mathbf{K}\rangle\|^{2}\geq 100\cdot\mathrm{e}^{1/2}\lambda_{r+1}^{2}\Bigr{]}\leq\frac{1}{100}.$
@@
(27) $\theta+\mu+\|\mathbf{E}\|\leq(1+10\mathrm{e}^{1/4})(\lambda_{r+1}+\mu)\leq 14(\lambda_{r+1}+\mu).$
@@
(28) $\left(\frac{1}{\theta+\mu}+\frac{1}{\lambda_{d}+\mu}\right)\leq\frac{2}{\lambda_{d}+\mu}.$
@@
The proof of Theorem 4.2 is now straightforward. To get a high probability bound, we use a simple boosting “trick”. We expect more refined results can be obtained directly.
Proof 4.9 (Proof of Theorem 4.2).
We first analyze the case $q=1$ (i.e. $\delta=1/100$). By Theorem 4.7, we are guaranteed that with probability at least $99/100$,
@@
When the event in Eq. 29 holds, Corollary 2.3 guarantees that,
(30) $\Bigg{\{}\forall\mu\geq 0:\frac{\|\mathbf{A}^{-1}\mathbf{b}-\mathsf{pcg}_{t}(\mu)\|_{\mathbf{A}_{\mu}}}{\|\mathbf{A}^{-1}\mathbf{b}\|_{\mathbf{A}_{\mu}}}\leq 2\exp\left(-\frac{2t}{\sqrt{28\kappa_{r+1}(\mu)}}\right)\Bigg{\}}.$
The result follows from Theorem 3.4 and that $\sqrt{28}/2<3$.
@@
(31) $\mathcal{K}_{t}(\mathbf{A},[\mathbf{b}~\mathbf{\Omega}_{i}])\subseteq\mathcal{K}_{t}(\mathbf{A},\mathbf{\Omega})$
@@ 4.2 Improved matrix-vector product complexity
In this section, we prove Theorem 4.3. The general approach is identical to the proof of Theorem 4.2, but we use a different bound for Nyström low-rank approximation.
@@
Note that Theorem 4.10 is not for Nyström low-rank approximation. However, Nyström low-rank approximation is better than projection-based low-rank approximation [61], which allows us to prove Theorem 4.3.
Proof 4.11 (Proof of Theorem 4.3).
We set $\varepsilon=10\mathrm{e}^{1/4}-1$. Since $\operatorname{range}(\mathbf{Q})\subseteq\operatorname{range}(\mathbf{K}_{s})$, the standard monotonicity property of the Nyström approximation implies $\|\mathbf{A}-\mathbf{A}\langle\mathbf{K}_{s}\rangle\|\leq\|\mathbf{A}-\mathbf{A}\langle\mathbf{Q}\rangle\|$. Next, [61, Lemma 5.2] asserts that for any matrix $\mathbf{Q}$ with orthonormal columns, $\|\mathbf{A}-\mathbf{A}\langle\mathbf{Q}\rangle\|\leq\|\mathbf{A}-\mathbf{Q}\mathbf{Q}^{\mathsf{T}}\mathbf{A}\|$. Thus, by Theorem 4.10 (with $\varepsilon$ set to an appropriate constant), the given condition on $s$ ensures that, with probability at least $1-\delta$, $\|\mathbf{A}-\mathbf{A}\langle\mathbf{K}_{s}\rangle\|<10\mathrm{e}^{1/4}\lambda_{r+1}$. The proof then follows similarly to the proof of Theorem 4.2.
5 Discussion and comparison with past work
The block-CG algorithm is widely used to solve linear systems with multiple right hand sides, and in such settings working over the block Krylov subspace is often advantageous compared to working over the individual Krylov subspaces [48, 41, 22, 4]. Block-CG is also used to solve single linear systems. For instance, so-called enlarged KSMs split the right hand size vector $\mathbf{b}$ into multiple vectors using a domain decomposition approach, and then apply block-CG to this collection of vectors [30, 31, 32]. Augmenting a KSM with random vectors was shown to be beneficial in numerical experiments appearing in concurrent work [66, Appendix B].
While we are unaware of any bounds similar to Theorem 4.3, bounds similar to Theorem 4.2 are known. In particular, [48, §4] derives a convergence bound for block-CG terms of $\sqrt{\lambda_{1}/\lambda_{d-m+1}}$; i.e. the condition number of $\mathbf{A}$ with the bottom $m$ eigenvalues removed. [^5] Note that the ordering of the eigenvalues in [48] is reversed from the present paper. At a high level, [48] obtains this bound by showing that the block Krylov subspace contains vectors which can approximately annihilate a subset of the eigenvalues of $\mathbf{A}$. Thus, in principle, other sets of eigenvalues (besides the bottom ones) can be eliminated. The bounds of [48] work for any starting block (but as a result, are more technical, depending on how “good” the starting block is). It would not be surprising such quantities can be bounded when the starting block is Gaussian.
KSMs are widely used for both low-rank approximation and problems involving matrix-functions, but their behavior on these problems is fundamentally different. As such, it would be interesting to try to avoid relying on a black-box reduction to low-rank approximation. For example, while low-rank approximation is complicated by the presence of repeated eigenvalues, solving linear systems is not. The approach of [48] is closely related to the approaches used in [61, 45] to prove bounds for low-rank approximation, so it is conceivable that, by taking the approach of [48], but using modern tools from randomized numerical linear algebra, one may be able to directly derive better bounds for randomized versions of block-CG.
More broadly, there are many randomized algorithms for linear systems and regression. For instance, a recent line of work shows that certain stochastic iterative methods enjoy an implicit preconditioning-like effect [14, 39, 17, 18, 16], with convergence independent of the top eigenvalues. These algorithms have runtime guarantees better than KSMs in some settings [16]. In addition, for tall least squares problems, the sketch-and-precondition framework for building a preconditioner for KSMs is extremely effective [54, 40], resulting in algorithms with optimal theoretical complexity [13, 11]. A number of recent works in randomized numerical linear algebra [10, 61, 44, 8] have made use of various nestedness properties of block-KSMs similar in flavor to Theorem 3.2 in order to obtain more efficient algorithms.
@@ 6 Gaussian sampling
(32) $\bm{\mu}+\mathbf{A}^{1/2}\mathbf{b}\sim\mathcal{N}(\bm{\mu},\mathbf{A}).$
In other words, in order to sample Gaussian vectors with covariance $\mathbf{A}$, it suffices to apply the matrix-square root of $\mathbf{A}$ to a standard Gaussian vector and then shift this result by $\bm{\mu}$. The most computationally difficult part of this is applying the square root of $\mathbf{A}$ to $\mathbf{b}$.
Recall the relation:
Eq. 4 $\mathbf{A}^{1/2}\mathbf{b}=\frac{2}{\pi}\int_{0}^{\infty}\mathbf{A}(\mathbf{A}+z^{2}\mathbf{I})^{-1}\mathbf{b}\,\mathrm{d}z.$
@@
This and closely related methods appear throughout the literature [12, 52, 7].
Often we wish to sample multiple vectors from $\mathcal{N}(\bm{\mu},\mathbf{A})$, and so we might use a block-variant of Definition 6.1.
@@
Both algorithms can be efficiently implemented using the (block)-Lanczos algorithms; see e.g. [7].
We can use the bounds for augmented block-CG from Section 4 to derive bounds for the block-Lanczos square root algorithm. In particular, using Theorem 4.2, we prove:
@@
Similarly, using Theorem 4.3, we prove:
@@
Proof 6.5 (Proof of Theorems 6.3 and 6.4).
Our proof is optimized for simplicity rather than sharpness. For notational convenience, let $\mathbf{b}=\mathbf{b}^{(1)}$ and $\mathbf{\Omega}=[\mathbf{b}^{(2)},\ldots,\mathbf{b}^{(m)}]$. As noted in Section 6, $\mathbf{\Omega}\in\mathbb{R}^{d\times(m-1)}$ is independent of $\mathbf{b}$. Consider the event
@@
(34) $\varepsilon_{t}^{\prime}(\mu)=2\exp\left(-\frac{t-s}{3\sqrt{\kappa_{r+1}(\mu)}}\right).$
To prove Theorem 6.3, note that the choice of $r$ ensures that $m-1\geq(r+2)q$ for $q=\lceil\log(m/\delta)/\log(100)\rceil$. Therefore, with $s=2+\log(d)/2$, Theorem 4.2 guarantees that Eq. 33 holds with probability at least $1-\delta/m$. Likewise, to prove Theorem 6.4, note that the Theorem 4.3 guarantees that Eq. 33 holds with probability at least $1-\delta/m$ (where we have used that $m\leq d$ so that $\log(dm/\delta)=O(\log(d/\delta))$).
From this point on we condition on Eq. 33, and the two proofs are identitcal. Applying standard norm inequalities we obtain a bound
(35) $\|\mathbf{A}\mathbf{A}_{\mu}^{-1}\mathbf{b}-\mathbf{A}\mathsf{bcg}_{t}^{(1)}(\mu)\|$ $=\|\mathbf{A}\mathbf{A}_{\mu}^{-1/2}\mathbf{A}_{\mu}^{1/2}(\mathbf{A}_{\mu}^{-1}\mathbf{b}-\mathsf{bcg}_{t}^{(1)}(\mu))\|$
(36) $\leq\|\mathbf{A}\mathbf{A}_{\mu}^{-1/2}\|\|\mathbf{A}_{\mu}^{-1}\mathbf{b}-\mathsf{bcg}_{t}^{(1)}(\mu)\|_{\mathbf{A}_{\mu}}.$
Under the assumption the event in Eq. 33 holds, we bound
(37) $\|\mathbf{A}_{\mu}^{-1}\mathbf{b}-\mathsf{bcg}_{t}^{(1)}(\mu)\|_{\mathbf{A}_{\mu}}\leq\varepsilon_{t}(\mu)\|\mathbf{A}_{\mu}^{-1}\mathbf{b}\|_{\mathbf{A}_{\mu}}\leq\varepsilon_{t}(\mu)\|\mathbf{A}_{\mu}^{-1/2}\|\|\mathbf{b}\|.$
@@
(38) $\|\mathbf{A}\mathbf{A}_{\mu}^{-1/2}\|=\max_{i}\frac{\lambda_{i}}{\sqrt{\lambda_{i}+\mu}}=\frac{\lambda_{1}}{\sqrt{\lambda_{1}+\mu}}$
@@
(39) $\|\mathbf{A}_{\mu}^{-1/2}\|=\max_{i}\frac{1}{\sqrt{\lambda_{i}+\mu}}=\frac{1}{\sqrt{\lambda_{d}+\mu}}.$
@@
Therefore, applying the triangle inequality for integrals and Eq. 40 we compute a bound
(41) $\|\mathbf{A}^{1/2}\mathbf{b}-\mathsf{bsq}_{t}^{(1)}\|$ $=\frac{2}{\pi}\biggl{\|}\int_{0}^{\infty}\mathbf{A}(\mathbf{A}+z^{2}\mathbf{I})^{-1}\mathbf{b}-\mathbf{A}\mathsf{bcg}_{t}^{(1)}(z^{2})\mathrm{d}{z}\biggr{\|}$
(42) $\leq\frac{2}{\pi}\int_{0}^{\infty}\|\mathbf{A}(\mathbf{A}+z^{2}\mathbf{I})^{-1}\mathbf{b}-\mathbf{A}\mathsf{bcg}_{t}^{(1)}(z^{2})\|\mathrm{d}{z}$
@@
where $K(m):=\int_{0}^{\pi/2}(1-m\sin^{2}(z))^{-1/2}\mathrm{d}{z}$ is the complete elliptic integral of the first kind. Standard bounds on elliptic integrals (see Lemma C.1) guarantee that, for all $x>1$,
@@
Therefore, applying Eq. 45 with $x=\kappa=\lambda_{1}/\lambda_{d}$ to Eqs. 43 and 44 we get a bound
(46) $\|\mathbf{A}^{1/2}\mathbf{b}-\mathsf{bsq}_{t}^{(1)}\|\leq\frac{\sqrt{\lambda_{1}}}{2}\log\left(16\kappa\right)\varepsilon_{t}^{\prime}(0)\|\mathbf{b}\|.$
Since $\sqrt{\lambda_{1}}=\|\mathbf{A}^{1/2}\|$, this is the desired result for $\mathbf{b}^{(1)}$.
To obtain the final result, we observe that $\mathbf{b}^{(1)},\ldots,\mathbf{b}^{(m)}$ and $\mathsf{bsq}_{t}^{(1)},\ldots,\mathsf{bsq}_{t}^{(m)}$ are permutation invariant in distribution. Thus, the above result in fact applies to each $\mathbf{b}^{(i)}$ and $\mathsf{bsq}_{t}^{(i)}$ pair with probability at least $1-\delta/m$. Applying a union bound over these $m$ events gives the main result.
Theorems 6.3 and 6.4 gives bounds on the matrix-vector products required by the block-Lanczos method. For instance, Theorem 6.3 implies we can use roughly
(47) $m\Big{(}\log(d)+\sqrt{\kappa_{r+1}}\log\big{(}\log(\kappa)/\varepsilon\big{)}\Big{)}$
matrix-vector products, where $r=O(m/\log(m))$. In contrast, existing bounds for methods like the single-vector Lanczos square root method Definition 6.1 [9, 52] sample a single Gaussian vector with roughly $\sqrt{\kappa}\log(1/\varepsilon)$ matrix-vector products. Therefore, when $\lambda_{r+1}\ll\lambda_{1}$, the total number of matrix-vector products is reduced significantly by using the block method in Definition 6.2.

### v3

@@ 4 Probabilistic bounds
In this section we prove probabilistic bounds for augmented block-CG. In light of Theorem˜3.4, our strategy for deriving bounds for block-CG is simply to derive bounds for Nyström PCG. To do this, we follow the approach developed in [25] and use a deterministic bound for the condition number of the Nyström preconditioned system. Note that block-CG always performs at least as well as CG (in terms of matrix-loads), and hence is guaranteed to converge (in exact arithmetic).
The following is a minor generalization of Proposition 5.3 in [25] to allow arbitrary $\theta$. The proof is contained in Subsection˜C.1 for completeness.
@@
Let $\mathbf{P}_{\mu}$ be the Nyström preconditioner Equation˜8 corresponding to the Nyström approximation $\mathbf{A}\langle\mathbf{K}\rangle$ for any $\mathbf{K}$ and shift parameter $\theta\geq 0$. Then
@@
In particular, Proposition˜4.1 implies that if $\|\mathbf{A}-\mathbf{A}\langle\mathbf{K}\rangle\|\approx\lambda_{r+1}$ and $\theta\in[\lambda_{d},\lambda_{r+1}]$, then $\kappa(\mathbf{P}_{\mu}^{-1/2}\mathbf{A}_{\mu}\mathbf{P}_{\mu}^{-1/2})\approx\kappa_{r+1}(\mu)$.
@@
Our first result, which we prove in Subsection˜4.1, is a “numerical analysis” style bound based on the error gurantees for Nyström low-rank approximation in [68].
@@
$s\geq\min\left\{3+\frac{\log(d)}{2},{\frac{3}{2}}+\frac{1}{4}\log\left(4+\frac{4r}{p-1}\sum_{i>r}\frac{\lambda_{i}^{2}}{\lambda_{r+1}^{2}}\right)\right\}.$
@@
Note that $\mathbf{A}\langle\mathbf{K}_{s}\rangle$ as defined in Equation˜9 can have rank as large as $\ell s$, so we might hope that we can deflate roughly $\ell s$ eigenvalues. Recent work [50, 9] implies this is more-or-less the case. In Subsection˜4.2, we use this result to prove our second result, a “theoretical computer science” type bound.
@@
In contrast to Theorem˜4.2, which requires $\ell\geq r+2$, Theorem˜4.3 allows any choice of $\ell$. The bound reveals that we get exponential convergence at a rate roughly $1/\sqrt{\kappa_{r+1}(\mu)}$ after roughly $s\approx r/\ell$ matrix-loads (and hence $r\approx\ell s$ matrix-vector products). Note that the logarithmic dependence on eigenvalue gaps $\Delta$ is generally considered to be mild due to “smoothed-analysis” type effects of finite-precision arithmetic [50, 9].
@@
The fact that Theorems˜4.2 and 4.3 guarantee an accurate result for all $\mu\geq 0$ with high probability (as opposed to a single value of $\mu$) will be important in our applications. In particular, it allows guarantees for computing the whole ridge-regression regularization path and will be necessary in the analysis of our algorithm for sampling Gaussian vectors.
@@
Bounds based on condition numbers (such as Theorem˜4.2) are often pessimistic in practice. To determine how many iterations to run, it is more common to use some sort of a posteriori error estimate. For example, observing the residual $\|\mathbf{b}-\mathbf{A}_{\mu}\mathsf{bcg}_{t}^{(1)}(\mu)\|$ gives some indication of the quality of the solution. More advanced techniques can also be used [49].
@@ 4.1 Explicit bound
In this section, we prove Theorem˜4.2. We do not attempt to optimize constants, opting instead for simple arguments and clean theorem statements.
@@
Suppose $\mathbf{\Omega}\in\mathbb{R}^{d\times(r+p)}$ is a random Gaussian matrix and define $\mathbf{K}_{s}:=[\mathbf{\Omega}~\mathbf{A}\mathbf{\Omega}~\cdots~\mathbf{A}^{s-1}\mathbf{\Omega}]$. Let $\mathbf{P}_{\mu}$ be the Nyström preconditioner Equation˜8 corresponding to the Nyström approximation $\mathbf{A}\langle\mathbf{K}_{s}\rangle$ for any shift parameter $\theta\in[\lambda_{d},\lambda_{r+1}]$. Suppose $p\geq 2$ and
$s\geq\min\left\{3+\frac{\log(d)}{2},\frac{3}{2}+\frac{1}{4}\log\left(4+\frac{4r}{p-1}\sum_{i>r}\frac{\lambda_{i}^{2}}{\lambda_{r+1}^{2}}\right)\right\}.$
@@
The proof is a simple consequence of Theorems˜4.6 and 4.1. We begin by bounding $\|\mathbf{A}-\mathbf{A}\langle\mathbf{K}\rangle\|$. We first note that the minimum is always attained by the second term. Indeed, since $r\leq d$ and $\lambda_{i}/\lambda_{r+1}\leq 1$ for $i>r$, so since $p\geq 2$, we see $4r/(p-1)\sum_{i>r}\lambda_{i}^{2}/\lambda_{r+1}^{2}\leq 4d^{2}$. Now, using properties of the logarithm, $\log(4+4d^{2})=\log(1+d^{2})+\log(4)\leq\log(d^{2})+\log(2)+\log(4)=\log(8)+2\log(d)$. The claim then follows since $3/2+\log(8)/4\leq 3$.
With the condition on $s$, Theorem˜4.6 guarantees that
$\operatorname{\mathbb{E}}\Bigl[\|\mathbf{A}-\mathbf{A}\langle\mathbf{K}\rangle\|^{2}\Bigr]\leq\mathrm{e}^{1/2}\cdot\lambda_{r+1}^{2}.$
@@
$\operatorname{\mathbb{P}}\Bigl[\|\mathbf{A}-\mathbf{A}\langle\mathbf{K}\rangle\|^{2}\geq 100\cdot\mathrm{e}^{1/2}\lambda_{r+1}^{2}\Bigr]\leq\frac{1}{100}.$
@@
$\theta+\mu+\|\mathbf{E}\|\leq(1+10\mathrm{e}^{1/4})(\lambda_{r+1}+\mu)\leq 14(\lambda_{r+1}+\mu).$
@@
$\left(\frac{1}{\theta+\mu}+\frac{1}{\lambda_{d}+\mu}\right)\leq\frac{2}{\lambda_{d}+\mu}.$
@@
The proof of Theorem˜4.2 is now straightforward. To get a high probability bound, we use a simple boosting “trick”. We expect more refined results can be obtained directly.
Proof of Theorem˜4.2. We first analyze the case $q=1$ (i.e. $\delta=1/100$). By Theorem˜4.7, we are guaranteed that with probability at least $99/100$,
@@
When the event in Equation˜11 holds, Corollary˜2.3 guarantees that,
$\Bigg\{\forall\mu\geq 0:\frac{\|\mathbf{A}^{-1}\mathbf{b}-\mathsf{pcg}_{t}(\mu)\|_{\mathbf{A}_{\mu}}}{\|\mathbf{A}^{-1}\mathbf{b}\|_{\mathbf{A}_{\mu}}}\leq 2\exp\left(-\frac{2t}{\sqrt{28\kappa_{r+1}(\mu)}}\right)\Bigg\}.$
The result follows from Theorem˜3.4 and that $\sqrt{28}/2<3$.
@@
$\mathcal{K}_{t}(\mathbf{A},[\mathbf{b}~\mathbf{\Omega}_{i}])\subseteq\mathcal{K}_{t}(\mathbf{A},\mathbf{\Omega})$
@@ 4.2 Improved matrix-vector product complexity
In this section, we prove Theorem˜4.3. The general approach is identical to the proof of Theorem˜4.2, but we use a different bound for Nyström low-rank approximation.
@@
Note that ˜4.9 is not for Nyström low-rank approximation. However, Nyström low-rank approximation is better than projection-based low-rank approximation [68], which allows us to prove Theorem˜4.3. Proof of Theorem˜4.3. We set $\varepsilon=10\mathrm{e}^{1/4}-1$. Since $\operatorname{range}(\mathbf{Q})\subseteq\operatorname{range}(\mathbf{K}_{s})$, the standard monotonicity property of the Nyström approximation implies $\|\mathbf{A}-\mathbf{A}\langle\mathbf{K}_{s}\rangle\|\leq\|\mathbf{A}-\mathbf{A}\langle\mathbf{Q}\rangle\|$. Next, [68, Lemma 5.2] asserts that for any matrix $\mathbf{Q}$ with orthonormal columns, $\|\mathbf{A}-\mathbf{A}\langle\mathbf{Q}\rangle\|\leq\|\mathbf{A}-\mathbf{Q}\mathbf{Q}^{\mathsf{T}}\mathbf{A}\|$. Thus, by ˜4.9 (with $\varepsilon$ set to an appropriate constant), the given condition on $s$ ensures that, with probability at least $1-\delta$, $\|\mathbf{A}-\mathbf{A}\langle\mathbf{K}_{s}\rangle\|<10\mathrm{e}^{1/4}\lambda_{r+1}$. The proof then follows similarly to the proof of Theorem˜4.2.
Remark 4.10.
A number of algorithms for Gaussian process regression make use of block-KSMs to simultaneously solve a linear system and apply a matrix-function to a collection of Gaussian random vectors (for stochastic trace estimation) [30, 72]. Our bounds are relevant to this setting, and provide further theoretical justification to the use of block-KSMs in these settings.
@@ 5 Gaussian sampling
$\bm{\mu}+\mathbf{A}^{1/2}\mathbf{b}\sim\mathcal{N}(\bm{\mu},\mathbf{A}).$
In other words, in order to sample Gaussian vectors with covariance $\mathbf{A}$, it suffices to apply the matrix-square root of $\mathbf{A}$ to a standard Gaussian vector and then shift this result by $\bm{\mu}$. The most computationally difficult part of this is applying the square root of $\mathbf{A}$ to $\mathbf{b}$. Recall the relation:
$\mathbf{A}^{1/2}\mathbf{b}=\frac{2}{\pi}\int_{0}^{\infty}\mathbf{A}(\mathbf{A}+z^{2}\mathbf{I})^{-1}\mathbf{b}\,\mathrm{d}z.$
@@
This and closely related methods appear throughout the literature [13, 59, 8]. Often we wish to sample multiple vectors from $\mathcal{N}(\bm{\mu},\mathbf{A})$, and so we might use a block-variant of ˜5.1.
@@
Both algorithms can be efficiently implemented using the (block)-Lanczos algorithms; see e.g. [8]. We can use the bounds for augmented block-CG from Section˜4 to derive bounds for the block-Lanczos square root algorithm. In particular, using Theorem˜4.2, we prove:
@@
Similarly, using Theorem˜4.3, we prove:
@@
[…]
@@
$\varepsilon_{t}^{\prime}(\mu)=2\exp\left(-\frac{t-s}{3\sqrt{\kappa_{r+1}(\mu)}}\right).$
To prove ˜5.3, note that the choice of $r$ ensures that $m-1\geq(r+2)q$ for $q=\lceil\log(m/\delta)/\log(100)\rceil$. Therefore, with $s=2+\log(d)/2$, Theorem˜4.2 guarantees that ˜12 holds with probability at least $1-\delta/m$. Likewise, to prove ˜5.4, note that the Theorem˜4.3 guarantees that ˜12 holds with probability at least $1-\delta/m$ (where we have used that $m\leq d$ so that $\log(dm/\delta)=O(\log(d/\delta))$). From this point on we condition on ˜12, and the two proofs are identical. Applying standard norm inequalities we obtain a bound
$\|\mathbf{A}\mathbf{A}_{\mu}^{-1}\mathbf{b}-\mathbf{A}\mathsf{bcg}_{t}^{(1)}(\mu)\|$ $=\|\mathbf{A}\mathbf{A}_{\mu}^{-1/2}\mathbf{A}_{\mu}^{1/2}(\mathbf{A}_{\mu}^{-1}\mathbf{b}-\mathsf{bcg}_{t}^{(1)}(\mu))\|$
$\leq\|\mathbf{A}\mathbf{A}_{\mu}^{-1/2}\|\|\mathbf{A}_{\mu}^{-1}\mathbf{b}-\mathsf{bcg}_{t}^{(1)}(\mu)\|_{\mathbf{A}_{\mu}}.$
Under the assumption the event in ˜12 holds, we bound
$\|\mathbf{A}_{\mu}^{-1}\mathbf{b}-\mathsf{bcg}_{t}^{(1)}(\mu)\|_{\mathbf{A}_{\mu}}\leq\varepsilon_{t}(\mu)\|\mathbf{A}_{\mu}^{-1}\mathbf{b}\|_{\mathbf{A}_{\mu}}\leq\varepsilon_{t}(\mu)\|\mathbf{A}_{\mu}^{-1/2}\|\|\mathbf{b}\|.$
@@
$\|\mathbf{A}\mathbf{A}_{\mu}^{-1/2}\|=\max_{i}\frac{\lambda_{i}}{\sqrt{\lambda_{i}+\mu}}=\frac{\lambda_{1}}{\sqrt{\lambda_{1}+\mu}}$
@@
$\|\mathbf{A}_{\mu}^{-1/2}\|=\max_{i}\frac{1}{\sqrt{\lambda_{i}+\mu}}=\frac{1}{\sqrt{\lambda_{d}+\mu}}.$
@@
Therefore, applying the triangle inequality for integrals and ˜13 we compute a bound
$\|\mathbf{A}^{1/2}\mathbf{b}-\mathsf{bsq}_{t}^{(1)}\|$ $=\frac{2}{\pi}\biggl\|\int_{0}^{\infty}\mathbf{A}(\mathbf{A}+z^{2}\mathbf{I})^{-1}\mathbf{b}-\mathbf{A}\mathsf{bcg}_{t}^{(1)}(z^{2})\mathrm{d}{z}\biggr\|$
$\leq\frac{2}{\pi}\int_{0}^{\infty}\|\mathbf{A}(\mathbf{A}+z^{2}\mathbf{I})^{-1}\mathbf{b}-\mathbf{A}\mathsf{bcg}_{t}^{(1)}(z^{2})\|\mathrm{d}{z}$
@@
where $K(m):=\int_{0}^{\pi/2}(1-m\sin^{2}(z))^{-1/2}\mathrm{d}{z}$ is the complete elliptic integral of the first kind. Standard bounds on elliptic integrals (see LABEL:thm:elliptic_k_bd) guarantee that, for all $x>1$,
@@
[…]
$\|\mathbf{A}^{1/2}\mathbf{b}-\mathsf{bsq}_{t}^{(1)}\|\leq\frac{\sqrt{\lambda_{1}}}{2}\log\left(16\kappa\right)\varepsilon_{t}^{\prime}(0)\|\mathbf{b}\|.$
[…]
$m\Big(\log(d)+\sqrt{\kappa_{r+1}}\log\big(\log(\kappa)/\varepsilon\big)\Big)$
matrix-vector products, where $r=O(m/\log(m))$. In contrast, existing bounds for methods like the single-vector Lanczos square root method ˜5.1 [10, 59] sample a single Gaussian vector with roughly $\sqrt{\kappa}\log(1/\varepsilon)$ matrix-vector products. Therefore, when $\lambda_{r+1}\ll\lambda_{1}$, the total number of matrix-vector products is reduced significantly by using the block method in ˜5.2.